position of particle moving along x axis is given by x=3t-4t^2+t^3 where x is in meter and t in second find average velocity of particle in time interval t=2 to t=4 sec

Grade:12th pass

2 Answers

Arun
5 years ago
differentiate the equation w.r.t. time n put value of T.
 
hope it helps
 
Regards
Arun (askIITians forum expert)
Khimraj
5 years ago
x = t3-4t2+3t
v = dx/dt = 3t2 – 8t + 3
average velocity = (1/T)\int_{0}^{T}vdt
(1/2)\int_{2}^{4}(3t^{2}-8t+3)dt
(1/2)[(t^{3}-4t^{2}+3t)]_{2}^{4}
= 14/2 = 7m/s
Hope it clears.................